The following cash flows are equivalent at 8% for what value of Y?

End of Year 0 1 2 3 4 5
CF1, $ -$15,000 $5,000 $5,000 $5,000 $5,000 $5,000
CF2, $ Y? $8,000 $7,000 $6,000 $5,000 $4,000
Answer #1

Formula

As the two cash flows are equivalent, we can say

4963.550185 = y+25249.83784

or, y = 4963.550185-25249.83784 = -20286.28766